OpenStudy (anonymous):

exact value of cot x=1/sqrt3

OpenStudy (lgbasallote):

take the arc cotangent of both sides. that should give you a satisfying answer..

OpenStudy (lgbasallote):

or you can change it to tangent first \[\cot x = \frac{1}{\sqrt 3}\] \[\implies \tan x = \sqrt 3\] then take the arc tan

OpenStudy (lgbasallote):

do you see in your calculator something that looks like \(\tan^{-1}\)

OpenStudy (lgbasallote):

what kind of calculator are you using @_@

OpenStudy (lgbasallote):

use wolfram then http://www.wolframalpha.com/input/?i=arc+tan+%28sqrt+3%29

OpenStudy (lgbasallote):

oh..

OpenStudy (lgbasallote):

then you need to memorize the trig thingies

OpenStudy (lgbasallote):

so your problem is just the quadrant?

OpenStudy (lgbasallote):

im not following your line sorry